OpenStudy (anonymous):

MEDAL AND FAN If each foot of water in a lake screens out 60% of the light above, what percent of the light passes through 5 feet of water? a. 15% c. 1.024% b. 25% d. 3.725%

OpenStudy (anonymous):

@D3xt3R help again ;)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

if the lake screens out 60% of the light above, then 40% pass by it.\[P(h)=0.40^{h}\]\[P(5)=0.40^5\]The correct anser is:\[P(5)=0.01024=1.024\%\]

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

any logic behind getting that equation so easily??

jimthompson5910 (jim_thompson5910):

Each foot screens out 60% of the light. So 100% - 60% = 40% of the light gets through in the first foot. Through the second foot of water, 40% of the light that made it through then makes it through to the next layer. So 0.4*0.4 so far Third foot: 0.4*0.4*0.4 etc etc Through the 5th foot of water, we'll have 0.4*0.4*0.4*0.4*0.4 = 0.4^5 which is how D3xt3R got that equation

OpenStudy (anonymous):

thank u